How many liters of a 5% solution of hydrochloric acid must be mixed with 12 liters of a 1% solution to obtain a 4% solution?
Answer by stanbon(75887)

How many liters of a 5% solution of hydrochloric acid must be mixed with 12 liters of a 1% solution to obtain a 4% solution?
-----
Equation:
acid + acid = acid
0.05x + 0.01*12 = 0.04(x+12)
----
Multiply thru by 100 to get:
5x + 12 = 4x + 4*12
---
x = 3*12
x = 36 liters (amt. of 5% solution needed)
---
Cheers,
Stan H.
=========